From nch@roe.ac.uk Fri Jun 19 17:26:09 2009 Date: Fri, 19 Jun 2009 17:12:33 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 19/6/09 Minutes of WFAU VDFS Science Archive meeting: 19th June 2009 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, MAR, MSH, NJC, ETWS Apologies: JPE, JMS, AL, RGM, PMW, RPB, RSC NB: DoNM: 10am, Friday June 26th 2009 in the Vista Hut Actions discharged this week: ----------------------------- ACTION: NCH to run a standard science query sanity check UKIDSSDR5-BestDR7 (example Query #4.2.6 from the WSA paper) Discharged; everything checks out fine ACTION: MAR to wsa-announce availability of SDSS BestDR7. Discharged ACTION: ETWS to communicate with MJI/MR at CASU over transfer of the reprocessed GPS and UDS products for DR5 Discharged; see below ACTION: ETWS to send MSH the SDSS DR7 metadocs so that it can be published through the VO services Discharged; a minor issue concerning funny column types in the XML metadoc iis being resolved today. ACTION: MSH to switch off the secured access to UKIDSS DR3 as it is now out of it's proprietary period Discharged; UKIDSSDR3 is now available through the VO as well. Actions partly discharged but continuing: ----------------------------------------- ACTION: ETWS to parse that latest keyword info into the VSA metadata schema files. Continues; ETWS plans to shake down the parsing process using the latest stacked pawprints, then run it in anger on tile products. ACTION: NCH to draft an interface control document for VSA to ESO-SAF Drafted (before minutes typed, if not the meeting...) Actions carried forward from 5/06/09 meeting: ---------------------------------------------- None this week Specific points and new actions: -------------------------------- Project management: Nothing of note this week. WFCAM & VISTA updates: Nothing new this week. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: Somebody dug a hole in Leeds during the week and stuck their spade through the UKLight fibres, so connectivity was lost for a couple of days. Local Area Network upgrades tomorrow mean that archive services will be offline for a few hours. WSA Operations: ETWS noted: - Enhanced CU1 for the download of the reprocessed UDS and GPS data out of non-standard paths. - Finished update of LGRs list-driven photometry code and included all parameters from RGMs wishlist. - Transferred and ingested 09A May data. - Transferred the reprocessed GPS data. - Started catalogue ingests for 08B: 100% processed, 15% ingested so far. - Created new set of browser pages for newly released non-surveys. MAR noted that 08B (DR6) QC has started. Hardware: The public SSA server is now fixed after a week or so of downtime; direct comparison with the cluster deployment is now possible. Software: NJC noted: "I have been testing the cirdr code and comparing it to older versions and the data in the pipeline. The simple tests pass fine and I am now testing the indexing and the deblending fixes. I have also updated the RequiredStack and RequiredMosaic tables with the correct smoothing parameter for extraction. I have also been testing some changes to the schema, which allow for a more sensible rearrangement of the extractor wrapper. I have completed the first set of tests on this. I have also started looking into CU9, the list-driven photometry wrapper. Finally, I have got the referee's report back from the paper on archiving multi-epoch data. There are some minor corrections to be made." MAR noted that he's been working on enhancing the UI to prepare and warn users if their SQL is suspect (e.g. cartesian joins and similar denial-of-service screw-ups). NCH noted that select/execute permissions seem to be wrong on quite a few interface user logins / all public servers / all DBs for the functions/stored procedures, e.g. hatshepsut.PSSA.dbo.fHtmLookUp(), amenhotep.UKIDSSDR3PLUS.dbo.fHtmLookUp() etc. so an audit might be a good idea... Survey Data Release: Nothing new to note. Non-survey Data Release: Around 40 new and/or updated non-survey DBs have been released to users over the last week or so. Astrogrid deployment & Data Analysis services: Following ETWS' review of the list-driven photometry codes, final testing is being done by RGMcM (who requested some enhancements). Then we roll-out the service to all (probly). Miscellaneous: Ken Chambers and Jim Heasley (PanSTARRS) are visiting the IfA next Wednesday. RGM has sent around a call for input to the latest NeSC them of Next Generation Sky Surveys and IT challenges: http://www.nesc.ac.uk/esi/themes/theme_11/ and also http://wiki.esi.ac.uk/NGSS_Kick-Off_Inputs Finally, early doors tonight at The Guilford Arms for Kona's leaving do. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================